Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Ole Hansen
Log entry time 23:58:44 on March 02, 2011
Entry number 349753
keyword=Swing Shift Summary
Crew:
Ole Hansen (SL)
Whitney Armstrong (TO)
Narbe Kalantarians (3rd)
Summary:
Finished data taking for N-delta kinematics #2. Changed to kinematics #3,
set up prescales, and resumed production data taking.
Timeline:
- 16:00 Production on kinematics #2 in progress at beginning of shift
- 20:20 Reached charge goal of 810 mC. Asking for access to move LHRS
from 12.5° to 38.5°.
- 21:10 Access and spectrometer move completed. MCC (Jay) is finishing a
A-line test plan, so we wait for beam.
- 21:45 Beam is available again. Moving target to empty position for
beam recovery procedure.
- 22:00 Raster checks. BeO pointing run.
- 22:30 Various test runs to establish production beam current and prescale
factors.
- 23:05 Beginning kinematics #3 production with run 2217.
- 23:45 Mystery HV alarm from LeCroy beamline crate. All channels
show green. We turn the entire crate off and back on, and the alarm clears.
Should not have affected the run in progress (2219),
but one never knows what's plugged in there. Starting new run, 2220.
- 0:00 Shift ends with production for kinematics #3 in progress.
63 mC accumulated so far (runs 2217-2219)
Run list:
- 2194: production kinematics #2
- 2195-2196: junk, DAQ problems
- 2197-2206: production kinematics #2
- 2207-2208: cosmics
- 2209: Spot++, BeO target
- 2210: BeO pointing run
- 2211-2216: junk, beam current & prescale adjustments
- 2217-2220 (in progress): production, kinematics #3